5X^3+40Y^6 factor completely
Answer by stanbon(75887)

5X^3+40Y^6 factor completely
------
5(x^3 + 8y^6)
---
= 5(x^3 + (2y^2)^3)
----
= 5(x + 2y)(x^2-x(2y)+(2y)^2)
----
= 5(x+2y)(x^2-2xy+4y^2)
=============================
cheers,
Stan H.